Greenhab Report – February 13th

Crew 205 Green Hab Report – 13-FEB-2019

GreenHab Officer: Nathan Hadland

Environmental control: Heating and Cooling w/ ambient air (0 hrs)

Shade Cloth (40%)

Average temperatures:

Low: 17.4°C

High: 27.3°C

Hours of supplemental light: 5 Hours

Daily water usage for crops: 12 gallons

Water in Blue Tank: ~ 264.5 gallons

Time(s) of watering for crops:


1) 09.00 AM 21.4° C / 28 %
2) 12.00 PM 22.2° C / 24 %
3) 15:30 PM 21.9° C / 27 %
4) 18:15 PM 25.5° C / 33 %

Changes to crops: New seedlings sprouting around the large tomato plant in the grow bed.

Narrative: We continued regolith sample collection in Candor Chasma and will continue characterization throughout the evening using the dissection and optical microscopes. We are also performing spectral analysis using a spectrometer for preliminary elemental characterization.

The A. Thaliana has not germinated yet on the agar germination plates, so we are considering using a damp paper towel germination as a backup. The GreenHab dropped in temperature today due to persistent cloud cover. However, the heater mitigated the temperature fluctuations.

The red leaf lettuce us not showing any pigmentation, which is a possible indicator of plant stress. Additionally, the peas still have yellow pigmentation, which leads us to believe there are nutrient and nitrogen deficiency in those pots. We will continue to monitor these experiments.

We are still waiting for the arrival of seeds.

Harvest: None.

Support/supplies needed: None

Copyright © The Mars Society. All rights reserved. | Main Site